aliensbrain
  • Home
  • Study
  • Quizzes
  • 🎤AI Practicefree
  • Notebooks
  • Community
  • Sign in
  • Computer Science and Information Technology Fundamentals
  • A fair dice is rolled twice. The probability that an odd ...
Multiple choice

A fair dice is rolled twice. The probability that an odd number will follow an even number is

  1. 1/2

  2. 1/3

  3. 1

  4. 1/4

  5. None of these

Reveal answer Fill a bubble to check yourself
D Correct answer
Explanation

Required probability= 1/2 * 1/2 = 1/4

Keep practicing — related questions

  • A fair dice is tossed two times. The probability that the second toss results in a value that is higher tha...
  • Which of the following applications of RAID technology takes data that needs to be stored, and distributes ...
  • Which of the following organisms have the maximum number of chromosomes?
  • Which of the following UNIX commands counts the number of lines, words, and characters in a file?
  • Which of the following has been characterized by the number crunching algorithms?
  • Directions: In the following question, a statement is followed by two conclusions. Which of the following c...
  • Directions: In the following question, a question word is followed by four illustrations. Choose the option...
  • Direction: In the following question, a question word is followed by four illustrations. Choose the option ...
Play the full quiz 🎤 Practise this topic out loud

Practice this topic

  • Probability (337 questions)
Advertisement
© Aliensbrain | all rights reserved
  • About
  • Contact
  • Terms and Condition
  • Privacy Policy